# Evaluating the effects of Data Sparsity on the Link-level Bicycling Volume Estimation: A Graph Convolutional Neural Network Approach

### v1

@@ Abstract
Accurate bicycling volume estimation is crucial for making informed decisions about future investments in bicycling infrastructure. Traditional link-level volume estimation models are effective for motorised traffic but face significant challenges when applied to the bicycling context because of sparse data and the intricate nature of bicycling mobility patterns. To the best of our knowledge, we present the first study to utilize a Graph Convolutional Network (GCN) architecture to model link-level bicycling volumes. We estimate the Annual Average Daily Bicycle (AADB) counts across the City of Melbourne, Australia using Strava Metro bicycling count data. To evaluate the effectiveness of the GCN model, we benchmark it against traditional machine learning models, such as linear regression, support vector machines, and random forest. Our results show that the GCN model performs better than these traditional models in predicting AADB counts, demonstrating its ability to capture the spatial dependencies inherent in bicycle traffic data. We further investigate how varying levels of data sparsity affect performance of the GCN architecture. The GCN architecture performs well and better up to 80% sparsity level, but its limitations become apparent as the data sparsity increases further, emphasizing the need for further research on handling extreme data sparsity in bicycling volume estimation. Our findings offer valuable insights for city planners aiming to improve bicycling infrastructure and promote sustainable transportation.
Keywords: Intelligent Transportation Systems, Graph Convolutional Networks (GCN), Graph Neural Networks, Machine Learning, Data Sparsity, Strava Metro Data Analysis, Bicycling Volume Estimation, Sustainable Urban Transportation
@@ 1 Introduction
Although GNNs have demonstrated state-of-the-art performance in predicting motor vehicle volumes (Li et al., , 2017; Yu et al., , 2017; Zhang et al., , 2018; Zhao et al., , 2019), they have not yet been applied to modeling bicycling volumes due to the limited availability of relevant data and associated challenges. This study aims to address this gap by developing a robust GCN architecture for link-level bicycling volume modeling in the City of Melbourne and evaluate the effects of data sparsity on model performance using Strava Metro Data (Lee and Sener, , 2021). Our primary objective is to develop a GCN architecture that maximizes the predictive accuracy of the Annual Average Daily Bicycle (AADB) count, a critical metric used by transportation planners to understand infrastructure demand. To validate the performance of our proposed GCN model, we benchmark it against traditional machine learning models, including linear regression, support vector machines, and random forest. Additionally, we systematically introduce varying degrees of sparsity into the bicycle network to explore the impact of sparsity on model performance. Our findings contribute to the advancement of predictive modeling techniques in bicycling volume estimation and offer insights for city planners and policymakers aiming to enhance bicycling infrastructure and promote sustainable transportation choices.
@@ 3 Datasets used
However, the City of Melbourne currently has only 141 permanent bicycle count locations which represent only 1.1% of the total network links. This limited coverage highlights the significant data sparsity in the available ground-truth observations, necessitating the need for advanced techniques to estimate bicycling volumes effectively across the entire network. To achieve a comprehensive understanding of the bicycling network within the City of Melbourne, we use a combination of infrastructure data (from OpenStreetMap) and crowd-sourced bicycle volume data (from Strava Metro).
@@ 3.1 OSM Network Data
OSM (OpenStreetMap contributors, , 2022; Haklay and Weber, , 2008) is an open-source data that serves as a foundational dataset for our study, providing a detailed and comprehensive representation of the transportation network within the study area. With our specific focus on bicycle-related infrastructure, we extract the road/path network data from OSM which includes the topology and geometry of the road network, as well as associated features such as the presence and location of dedicated bicycle infrastructure such as bike lanes, shared road spaces. This data allows us to construct the graph representation of the network, where nodes represent intersections or road segments, and edges represent the connections between them. By integrating this network data with our bicycle count datasets, we ensure a comprehensive approach to modeling bicycling volumes, thereby improving the accuracy of our predictions and enabling a more informed analysis of infrastructure needs across the city.
@@ 3.2 Strava Metro Data
Unlike traditional count data, which is limited to only 141 counters in the City of Melbourne, Strava Metro offers bicycling counts across all 15,933 links in the network. This extensive coverage enables us to simulate various levels of data sparsity by systematically reducing the available count data and examining its impact on the performance of our Graph Convolutional Network (GCN) model. Strava data thus provides a robust foundation for training the model on a fully populated network and evaluating its performance under different levels of sparsity, reflecting real-world conditions of data sparsity.
However, it is important to recognize the limitations of Strava Metro data. It tends to over-represent specific cyclist groups, such as recreational riders and fitness enthusiasts, which can introduce bias by emphasizing well-trafficked routes and underreporting quieter or less-traveled areas. This bias may skew the overall representation of cycling behavior (Nelson et al., , 2021; Venter et al., , 2023). Despite these limitations, the comprehensive nature of Strava Metro data remains a valuable asset for understanding how data sparsity affects model performance in link-level bicycling volume estimation.
@@ 4.1 Data preparation and processing
As the first step, we extracted the road/path network of the City of Melbourne from OpenStreetMap (OSM) using the OSMnx (Boeing, , 2017) library in Python. The OSM dataset provides unique OSM IDs for each feature, distinguishing nodes, ways, and relations, along with detailed attributes for each road/path, such as surface type, and speed limit. To ensure comprehensive coverage of all roads/paths where bicyclists are permitted to ride, we extracted relevant OSM data using OSMnx by employing a robust combination of tags and corresponding values. Subsequently, we utilized a bike infrastructure classification system (Sustainable Mobility and Safety Research Group, Monash University, , 2023) and bike level of traffic stress (LTS) model (Sustainable Mobility and Safety Research Group, Monash University, , 2024) bespoke to Greater Melbourne to appropriately classify the types of infrastructure and bicycle level of traffic stress for each street segment in the network, respectively.
@@
where $AADB_{i}$, $bicycle\_trip\_count_{ij}$ and $n_{i}$ represent the annual average daily bicycle count, count for the $j^{th}$ observation, total number of observations for the $i^{th}$ OSM ID respectively and $\lceil$.$\rceil$ . ⌉ denotes the ceiling function, rounding up to the nearest integer. This approach enables the comparison of bicycling volumes across various locations and also normalizes the temporal variations by smoothing out fluctuations caused by seasonal effects (e.g., higher bicycle usage in warmer months) and anomalies (e.g., sudden drops in counts due to extreme weather or special events). By considering data across the entire year, AADB provides a consistent measure that is less affected by short-term variations. As derived from Strava data, we have denoted it as Strava AADB.
To create a comprehensive dataset for modeling link-level bicycling volumes, we integrate the OSM data with the Strava, covering 15,933 links (road/paths) across the City of Melbourne and Strava’s Annual Average Daily Bicycle (AADB) count is used as the target parameter for training the Graph Convolutional Network (GCN). While Strava provides count data for all links, the primary focus of this study is to evaluate the GCN model’s performance under varying conditions of data sparsity. In real-world scenarios, comprehensive bicycle count data, such as that provided by Strava, is rarely available across an entire network. Many cities have limited sensor coverage, resulting in sparse and incomplete data. To simulate these real-world conditions, we leverage Strava’s full dataset by progressively reducing the available count data, introducing different levels of sparsity. This approach enables us to assess the robustness of the GCN model as it encounters fewer data points, mimicking the challenges faced by urban planners in cities with limited bicycle count infrastructure. This evaluation helps to determine how well the GCN can generalize and make accurate predictions when operating with extreme level of data sparsity, a common issue in real-world urban mobility contexts.
@@
Before configuring the input layer for the Graph Convolutional Network (GCN) architecture, we conducted several data enrichment steps to ensure the dataset was well-prepared for modeling bicycling volumes. We primarily used OpenStreetMap (OSM) data to capture the road network and bicycle infrastructure within the City of Melbourne. To enhance the robustness of the data for bicycling volume estimation, we supplemented missing or inaccurately recorded attributes in OSM, such as slope and speed limits respectively, by spatially merging the OSM network with additional datasets which are as follows - (Department of Energy, Environment and Climate Action Victoria, , 2021) (slope data) and (Department of Transport and Planning Victoria, , 2014) (speed limit data), ensuring a more comprehensive representation of key features relevant to bicycling volumes on road segments.
@@
Additionally, missing data in categorical variables such as road/path surface types and infrastructure classifications were addressed by imputing values using the mode while continuous variables such as segment length and speed limit were imputed using the mean. Given the diverse nature of the data, Min-Max scaling was applied to continuous predictors to normalize their values and ensure they were on a comparable scale. Categorical variables, which included classifications of road types and bicycle facilities, were converted into numerical format through one-hot encoding, enabling the GCN to process these features effectively. The Strava AADB count data exhibited significant positive skewness, meaning a majority of road segments had low bicycle volumes, while a small number of segments recorded disproportionately higher counts (as shown in Figure 1). Such skewness could lead to biased model predictions and hinder the model’s ability to generalize across different data segments. To avoid this, we applied several data transformations such as Log, Box-Cox, Square Root, Quantile, and Yeo-Johnson methods, aiming to address the skewness, normalise the distribution and improve model performance. Each transformation was evaluated for its ability to normalize the distribution and reduce skewness. The results in the Table 1 indicated that while all transformations improved the distribution’s symmetry, the Box-Cox and Yeo-Johnson transformations were particularly successful in achieving a more balanced distribution. Given its performance, we select the Box-Cox transformed Strava AADB count as the target variable for further use in the Graph Convolutional Network (GCN) modeling task.
With these data preparation and preprocessing steps completed, the next step was to transform the data into graph structures that represented the spatial and infrastructure relationships within the bicycle network. The following section will delve into the graph preparation process and its critical role in enhancing the GCN model’s ability to capture complex patterns in bicycling activity.
@@ 4.2 Graph Creation for GCN
This section details the construction of a graph representation of the prepared bicycling volume dataset to effectively evaluate the impact of data sparsity on the prediction of link-level bicycling volumes using a Graph Convolutional Network (GCN). The graph creation process leverages the combined data from OpenStreetMap (OSM) and Strava Metro data, ensuring a comprehensive and detailed representation of the bicycling network. The construction of the graph begins with the extraction of key components from the integrated dataset, including unique OpenStreetMap (OSM) IDs, which represent individual road or path segments, the connectivity between these segments (edges), and the corresponding Strava Annual Average Daily Bicycle (AADB) counts. These components are critical for accurately modeling the link-level bicycling volumes within the City of Melbourne. Each unique OSM ID in the dataset is treated as a node in the graph. Nodes represent individual road/path segments within the bicycling network. Node features are derived from the attributes associated with each OSM ID, such as road type, surface quality, speed limits, level of traffic stress, slope, and bike infrastructure classification. These features are essential for the GCN, as they provide the input data that the network uses to learn and make predictions. Edges in the graph are created by defining connections between nodes based on the physical connectivity of the road/paths. For instance, if two segments are directly connected in the real world (i.e., they share a common intersection), an edge is created between the corresponding nodes in the graph. The edge features capture the relational attributes between connected nodes, such as the distance between segments or shared infrastructure types, or similar infrastructure patterns. These features are critical for the GCN as they allow the model to incorporate the spatial relationships between different parts of the network. The graph creation process also involves generating permutations of OSM ID pairs to comprehensively map the network’s connectivity. This results in a detailed graph that encapsulates the complexity of the bicycling network, allowing the GCN to model both local and global dependencies effectively.
The final graph is structured in a format compatible with PyTorch Geometric, a library designed for deep learning to work on irregularly structured data such as graphs. The data structure includes tensors for node features, edge indices and edge attributes which together define the graph’s topology and its associated data. Strava AADB counts associated with the corresponding nodes is the the target values for the GCN. This allows the GCN to estimate the AADB for each road/path segments based on both its intrinsic properties (node features) and its relational context within the network (edge features). The transformation of the bicycling volume data into a graph structure enables the GCN to capture spatial dependencies and traffic flow patterns by aggregating information from neighboring nodes through convolution. This approach allows the model to estimate link-level volumes based on both a node’s features and its connections within the network. The graph format also facilitates the study of how data sparsity affects model performance.
@@ 4.4 Training Process
We divide the dataset into training, validation, and test sets with a split ratio of 80:5:15 to comprehensively evaluate model performance. The training dataset is used to learn model parameters, while the validation set is employed to tune hyperparameters and prevent overfitting. We have set 2500 epochs for model training with early stopping, which halts the training if the validation loss does not improve over a set number of epochs, preventing the model from overfitting and saving computational resources. This technique ensures that the model’s capacity does not exceed the complexity of the data, especially in sparse settings.

### v2

@@ Abstract
Accurate bicycling volume estimation is crucial for making informed decisions and planning about future investments in bicycling infrastructure. However, traditional link-level volume estimation models are effective for motorised traffic but face significant challenges when applied to the bicycling context because of sparse data and the intricate nature of bicycling mobility patterns. To the best of our knowledge, we present the first study to utilize a Graph Convolutional Network (GCN) architecture to model link-level bicycling volumes and systematically investigate the impact of varying levels of data sparsity (0%–99%) on model performance, simulating real-world scenarios. We have leveraged Strava Metro data as the primary source of bicycling counts across 15,933 road segments/links in the City of Melbourne, Australia. To evaluate the effectiveness of the GCN model, we benchmark it against traditional machine learning models, such as linear regression, support vector machines, and random forest. Our results show that the GCN model outperforms these traditional models in predicting Annual Average Daily Bicycle (AADB) counts, demonstrating its ability to capture the spatial dependencies inherent in bicycle traffic network. While GCN remains robust up to 80% sparsity, its performance declines sharply beyond this threshold, highlighting the challenges of extreme data sparsity. These findings underscore the potential of GCNs in enhancing bicycling volume estimation, while also emphasizing the need for further research on methods to improve model resilience under high-sparsity conditions. Our findings offer valuable insights for city planners aiming to improve bicycling infrastructure and promote sustainable transportation.
Keywords: Bicycling Volume Estimation, Graph Convolutional Networks (GCN), Data Sparsity, Urban Mobility, Strava Metro Data
@@ 1 Introduction
Although GNNs have demonstrated state-of-the-art performance in predicting motor vehicle volumes (Li et al., , 2017; Yu et al., , 2017; Zhang et al., , 2018; Zhao et al., , 2019), they have not yet been applied to modeling bicycling volumes due to the limited availability of relevant data and associated challenges. This study aims to address this gap by developing a robust GCN architecture for link-level bicycling volume modeling in the City of Melbourne and evaluate the effects of data sparsity on model performance using Strava Metro Data (Lee and Sener, , 2021). To validate the performance of our proposed GCN model, we benchmark it against traditional machine learning models, including linear regression, support vector machines, and random forest. We systematically introduce varying degrees of sparsity into the bicycle network to explore the impact of sparsity on model performance. Our findings contribute to the advancement of predictive modeling techniques in bicycling volume estimation and offer insights for city planners and policymakers aiming to enhance bicycling infrastructure and promote sustainable transportation choices.
@@ 3 Datasets used
However, ground-truth observations of bicycling counts cover only a small fraction of these links, underscoring the substantial data sparsity within the network (Costa and José, , 2020; Nelson et al., 2021a, ). To achieve a comprehensive understanding of the bicycling network within the City of Melbourne, we rely on infrastructure data from OpenStreetMap (OSM) and crowd-sourced bicycle volume data from Strava Metro in this study.
@@ 3.1 OSM Network Data
OSM (OpenStreetMap contributors, , 2022; Haklay and Weber, , 2008) is an open-source data that serves as a foundational dataset for our study, providing a detailed and comprehensive representation of the transportation network within the study area. With our specific focus on bicycle-related infrastructure, we extract the road/path network data from OSM which includes the topology and geometry of the road network, as well as associated features such as slope, the presence and location of dedicated bicycle infrastructure such as bike lanes, shared road spaces. This data allows us to construct the graph representation of the network, where nodes represent road segments/links, and edges represent the intersection between them. By integrating this network data with our bicycle count dataset, we adopt a graph-based modeling approach to achieve accurate bicycling volumes modeling and a more informed analysis of bicycle network across the city.
@@ 3.2 Strava Metro Data
Unlike traditional limited count data, Strava Metro offers bicycling counts across all 15,933 links in the network. This extensive coverage enables us to simulate various levels of data sparsity by systematically reducing the available count data and examining its impact on the performance of our Graph Convolutional Network (GCN) model. In other words, we train the GCN on a fully populated Strava dataset, then progressively remove portions of the data to replicate real-world data scenario, evaluating how the model’s performance degrades under more extreme sparsity conditions.
Despite its extensive coverage, it is important to recognize the limitations of Strava Metro data. It tends to over-represent specific cyclist groups, such as recreational riders and fitness enthusiasts, which can introduce bias by emphasizing well-trafficked routes and underreporting quieter or less-traveled areas. This bias may skew the overall representation of cycling behavior (Nelson et al., 2021b, ; Venter et al., , 2023). Nevertheless, the comprehensive nature of Strava Metro data serves as a valuable testbed for understanding how data sparsity affects link-level bicycling volume estimation under realistic conditions of incomplete observation or limited bicycle count data.
@@ 4.1 Data preparation and processing
4.1.1 Extracting the Bicycle Network from OSM
As the first step, we extracted the road/path network of the City of Melbourne from OpenStreetMap (OSM) using the OSMnx (Boeing, , 2017) library in Python. The OSM dataset provides unique OSM IDs for each feature, distinguishing nodes, ways, and relations, along with detailed attributes for each road/path, such as surface type, and speed limit. To ensure comprehensive coverage of all roads/paths where bicyclists are permitted to ride, we extracted relevant OSM data using OSMnx by employing a robust combination of tags and corresponding values. We then used a bike infrastructure classification system (Sustainable Mobility and Safety Research Group, Monash University, , 2023) and bike level of traffic stress (LTS) model (Sustainable Mobility and Safety Research Group, Monash University, , 2024) bespoke to Greater Melbourne to appropriately classify the types of infrastructure and bicycle level of traffic stress for each street segment in the network, respectively.
4.1.2 Calculating Strava AADB
@@
where $AADB_{i}$, $bicycle\_trip\_count_{ij}$ and $n_{i}$ represent the annual average daily bicycle count, count for the $j^{th}$ observation, total number of observations for the $i^{th}$ OSM ID respectively and $\lceil$.$\rceil$ . ⌉ denotes the ceiling function, rounding up to the nearest integer. Strava AADB offers a consistent measure for comparing bicycling volumes across diverse locations, smoothing out fluctuations caused by seasonal effects or short-term anomalies. Although it derives from Strava data, which can be biased toward specific cyclist groups (e.g., recreational or fitness-oriented riders), this measure captures a broad distribution of cycling activity across 15,933 segments in the City of Melbourne.
4.1.3 Integrating OSM and Strava Data
To form a comprehensive dataset for link-level bicycling volume modeling, we integrate the OSM-based road/path network with Strava AADB counts. Each segment is linked via a unique OSM identifier, ensuring consistency in how infrastructure attributes and cycling volumes are matched. This combined dataset serves as the foundation for training and evaluating the Graph Convolutional Network (GCN) approach used in our experiments.
While Strava provides counts for the entire network, real-world conditions often involve sparse or incomplete bicycle count data. To simulate these conditions, we progressively reduce the available Strava count data, creating varying levels of sparsity. This strategy allows us to analyze how model performance changes as data becomes more limited, reflecting challenges faced by urban planners in cities with few permanent counters or limited resources.
@@ 4.1.3 Integrating OSM and Strava Data
Before configuring the input for the Graph Convolutional Network (GCN) architecture, we conduct several data enrichment steps to ensure the dataset is well-prepared for modeling bicycling volumes. We supplement missing or inaccurately recorded attributes in OSM, such as slope and speed limits respectively, by spatially merging the OSM network with additional datasets which are as follows - (Department of Energy, Environment and Climate Action Victoria, , 2021) (slope data) and (Department of Transport and Planning Victoria, , 2014) (speed limit data), ensuring a more comprehensive representation of key features relevant to bicycling volumes on road segments.
4.1.4 Data Enrichment and Cleaning
Additionally, missing data in categorical variables such as road/path surface types and infrastructure classifications are imputed using the mode value, ensuring that the most common category filled any gaps while continuous variables such as segment length and speed limit are imputed using the mean, avoiding distortions caused by outliers. Given the diverse nature of the data, Min-Max scaling is applied to continuous predictors to normalize their values and ensure they are on a comparable scale. Categorical variables, which include classifications of road types and bicycle facilities, are converted into numerical format through one-hot encoding, enabling the GCN to process these features effectively.
4.1.5 Addressing Skewness in Strava AADB
@@
As shown in Figure 1, the Strava AADB distribution exhibits a right (positive) skew, with most road segments having relatively low bicycle volumes while a small portion have much higher counts. Such skewness can lead to biased model predictions and limit the model’s generalization. To mitigate this issue, we test multiple transformations, including Log, Box-Cox, Square Root, Quantile and Yeo-Johnson.
Table 1 shows that Box-Cox and Yeo-Johnson transformations yield the most substantial reduction in skewness. Given its consistency, we select the Box-Cox transformation for our target variable in the GCN modeling task.
@@
With these data preparation and preprocessing steps completed, the next step is to transform the data into graph structures that represent the spatial and infrastructure relationships within the bicycle network. The following section will delve into the graph preparation process and its critical role in enhancing the GCN model’s ability to capture complex patterns in bicycling activity.
Figure 2: Overview of methodology for link-level bicycling volume prediction using using a Graph Convolutional Network (GCN) approach. Data section shows integration of Strava Metro Data (crowd-sourced bicycle volume data) and OpenStreetMap (OSM) Data (road infrastructure data), linked via unique OSM IDs. The graph construction phase includes: (a) mapping the bicycle network within the City of Melbourne, (b) demonstrating the process of graph inversion on a small area, where road segments are converted into nodes and intersections into edges, and (c) forming the final node-centric graph representation. The model architecture section is representing the GCN Configuration G (discussed in 4.3.5 and 5.1)
@@ 4.2 Graph Creation for GCN
This section details the construction of a graph representation of the prepared bicycling volume dataset, enabling the evaluation of data sparsity effects on link-level bicycling volume prediction using a Graph Convolutional Network (GCN). The graph creation process integrates data from OpenStreetMap (OSM) and Strava Metro, ensuring a spatially coherent and data-rich representation of Melbourne’s bicycling network.
4.2.1 Inverting the Traditional OSM Representation
A critical consideration in constructing this graph is the inversion of the traditional OSM representation, which represents intersections as nodes and road segments as edges. Instead, in our graph, each road/path segment is represented as a node, and intersections are modeled as edges connecting these nodes (as shown in Figure 2). This node-centric approach allows for the direct assignment of bicycling volume estimates to individual road segments. Moreover, this inversion aligns with graph-based framework, enabling the GCN to effectively learn from both segment attributes and their spatial relationships.
4.2.2 Graph Construction Process
The graph construction involves following key steps:
  • Defining Nodes and Features: Each road or path segment, uniquely identified by its OSM ID, is represented as a node. Node features include attributes such as road type, surface quality, speed limit, level of traffic stress, slope, and bicycle infrastructure classification. Additionally, the Strava-derived Annual Average Daily Bicycle (AADB) count is assigned to each node as the target variable for prediction. These features provide rich contextual information, enabling the GCN to leverage both intrinsic road characteristics and network-wide relationships.
  • Establishing Edges and Relationships: Edges represent intersections or physical connections between road segments. If two segments share a common intersection in OSM, an edge is created between their corresponding nodes. Edge features capture spatial relationships such as shared infrastructure types, allowing the GCN to model dependencies within the network effectively.
  • Integration with PyTorch Geometric: The final graph is formatted for compatibility with PyTorch Geometric, comprising node feature tensors, edge indices, and edge attribute tensors. This structure enables the GCN to utilize both node attributes and edge relationships to predict bicycling volumes based on the intrinsic properties of the road segments and their network context.
The node-centric graph structure is therefore pivotal for several reasons - 1. It preserves spatial relationships between road segments, allowing the model to learn how bicycle volumes on connected segments influence each other, 2. it facilitates the use of libraries like PyTorch Geometric, optimized for node regression tasks. It allows for controlled masking of nodes to simulate varying levels of data sparsity, providing insights into model robustness under realistic data conditions. This graph structure serves as the foundation for the GCN, enabling it to aggregate information from connected nodes and learn complex patterns in bicycling activity. The subsequent section details the training process and how the GCN leverages this graph structure to learn and generalize across the network.
@@ 4.4 Training Process
Figure 3: Data split and sparsity simulation for the training process. The constructed graph consists of 15,933 nodes (road segments/links), divided into training (80% - 12,746 nodes), validation (5% - 797 nodes), and testing (15% - 2,390 nodes) sets. To simulate varying sparsity levels, the training data is progressively reduced from 0% sparsity (12,746 nodes) to 99% sparsity (127 nodes), enabling a systematic evaluation of model performance under different levels of data sparsity.
@@
We divide the dataset into training, validation, and test sets with a split ratio of 80:5:15 to comprehensively evaluate model performance as shown in Figure 3. The training dataset is used to learn model parameters, while the validation set is employed to tune hyperparameters and prevent overfitting. We have set 2500 epochs for model training with early stopping, which halts the training if the validation loss does not improve over a set number of epochs, preventing the model from overfitting and saving computational resources. This technique ensures that the model’s capacity does not exceed the complexity of the data, especially in sparse settings.